# DiRL: An Efficient Post-Training Framework for Diffusion Language Models

### v1

@@ 1 Introduction
Large Diffusion Language Models (dLLMs) have become a hot topic in NLP (Nie et al., 2025; Inception, 2025). The emergence of dLLMs such as LLaDA (Nie et al., 2025), Dream (Ye et al., 2025), Mercury (Inception, 2025) and Gemini Diffusion (Gemini, 2025), together with blockwise hybrids like SDAR (Cheng et al., 2025) that combine diffusion with traditional auto-regression (AR), confirms the scalability of this paradigm (Nie et al., 2024; Gong et al., 2024; Ni et al., 2025). Building on these results, a growing body of work now seeks to advance their performance in multi-modality (Yang et al., 2025; You et al., 2025), long-context modeling (Liu et al., 2025b; He et al., 2025) and inference efficiency (Wu et al., 2025b; a; Song et al., 2025). Although pre-training of dLLMs is now proven feasible, post-training of dLLMs, especially reinforcement learning (RL), remains underdeveloped, limiting dLLMs’ performance on math tasks and real-world deployment.
@@
The difficulty of dLLM post-training, especially RL, lies in the fact that the logits and derived policy cannot be computed exactly (Zhao et al., 2025a; Zhu et al., 2025b). In original fully bidirectional dLLMs, the generation order is unconstrained, making teacher-forcing style logit acquisition during SFT infeasible. Injecting uniform random noise into the output fails to reproduce the realistic inference step map, resulting in biased logits and a large mismatch between training and inference objectives (Zhao et al., 2025a; Wang et al., 2025b). Furthermore, in the RL stage, the absence of a KV cache further increases computational overhead (Liu et al., ; Ma et al., 2025; Song et al., 2025). Most existing dLLM-based RL efforts lack an inference-engine backend, efficient training–inference co-design, and fast rollouts with online model updates, which prevents the practical adoption of mature RL algorithms such as GRPO (DeepSeek-AI, 2024; Guo et al., 2025). Blockwise dLLMs partially alleviate these issues by restricting generation within blocks, enabling exact logit computation through blockwise forward passes (Cheng et al., 2025; Wang et al., 2025b). However, they do not fully resolve the train–inference mismatch in post-training, nor do they address the efficiency and algorithmic challenges of dLLM RL. How to achieve consistent training and inference while enabling scalable RL for dLLMs remains underexplored. To fill this gap, we introduce our dLLM RL algorithm DiPO, together with the training framework DiRL, which enforces training–inference consistency and enables efficient rollouts and policy optimization for blockwise dLLMs. We further present the state-of-the-art dLLM DiRL-8B-Instruct, as shown in Figure 2, Figure 2 and Figure 3.
@@
At the algorithm level, DiPO leverages the good property of blockwise dLLM to achieve the first unbiased GRPO implementation for dLLMs through efficient, unbiased logit computation. At the framework level, DiRL supports the two-stage (SFT-RL) post-training of dLLMs, aligning training and inference objectives while surpassing existing methods in efficiency as shown in Figure 2. Concretely, we integrate the efficient inference property of blockwise dLLMs with the efficient FlexAttention interface (Dong et al., 2024) and LMDeploy framework (InternLM, 2023) to enable fast rollout and online model updates in the API server. At the model level, based on high-quality math datasets, we train DiRL-8B-Instruct from SDAR-8B-Chat and achieve best performance on math tasks in dLLMs, even outperform Qwen2.5 Series (Qwen et al., 2024), the widely-acknowledged larger AR model, in AIME24, AIME25 (MAA, 2024; 2025) and OlympiadBench (He et al., 2024) as shown in Figure 2. Our contributions can be summarized as follows.
@@ 2.1 Block Diffusion Language Model
Blockwise Diffusion Language Models, as exemplified by BD3-LMs (Arriola et al., 2025) and SDAR (Cheng et al., 2025), primarily unify the global sequential dependency of AR models with the local parallel generation capability of dLLMs through a Semi-Autoregressive generation paradigm.
@@
Thanks to the combination of intra-block AR and inner-block diffusion, the KV cache mechanism can be applied to blockwise dLLMs, thus enhancing the computational efficiency of traditional dLLMs and simplifying the likelihood computation (Cheng et al., 2025; Wang et al., 2025a).
@@ 2.2 RL and its application in dLLM
Thanks to the introduction of blockwise dLLMs (Cheng et al., 2025; Wang et al., 2025a), the computational overhead of log-probability is greatly reduced. Building on this, Wang et al. (2025b) proposes TraceRL, which leverages the algorithmic efficiency of blockwise dLLMs to enable faster unbiased logit computation. However, due to the lack of attention acceleration during training and the absence of tight integration between training and inference frameworks, TraceRL still leaves significant room for engineering optimization. Moreover, it does not implement more fine-grained RL algorithms such as GRPO. To address these limitations, we introduce our efficient training framework DiRL and DiPO, the first unbiased implementation of GRPO for dLLMs.
@@ 3.1 SFT stage
We train on the OpenR1-Math dataset (OpenR1, 2025) distilled from GLM-4.6 (Zeng et al., 2025a; zai-org, 2025). We chose GLM-4.6 because currently, it is almost the best-performing open-source non-reasoning LLM on math tasks and yields reasoning trajectories of manageable length. As long-reasoning evaluation for diffusion models is still scarce, we cap the reasoning length at 8k token length, which is already the longest inference length reported for dLLMs.
@@ 3.2 RL stage
Moreover, we integrate the DAPO(Yu et al., 2025) algorithm, which uses the token-level policy gradient:
@@
We train on the Big-Math dataset (Albalak et al., 2025), a large-scale, high-quality math dataset verfied by math-verify, and specially designed for RL.
@@ 4 Engineering Optimization
As shown in Figure 2, the DiRL framework incorporates efficiency designs at both the model and system levels. At the model level, blockwise dLLM (Cheng et al., 2025) with FlexAttention (Dong et al., 2024) compute unbiased logits efficiently, and at the framework level, tight integration of the LMDeploy inference engine (InternLM, 2023) with the open-source distributed training library LLaMA-Factory (Zheng et al., 2024) enables online model updates for highly efficient training.
@@ 4.1 Blockwise dLLM with FlexAttention
At the model level, as mentioned earlier, we use blockwise dLLM, which is more efficient for logic computation than the original dLLMs in post-training. This is because it maintains an AR paradigm between blocks, performing denoising only within each block. Thanks to this design, as Figure 4(a) shows, TraceRL (Wang et al., 2025b) can parallelize denoising across blocks by repeating the output twice and altering the attention mask, thus enabling faster SFT training.
However, the dominant FlashAttention interface (Dao et al., 2022; Dao, 2023) cannot handle the complex attention masks required by blocksize dLLMs, limiting the practical training throughput. Fortunately, PyTorch’s FlexAttention (Dong et al., 2024) fills this gap with an efficient operator that accepts fine-grained masks. DiRL does not differentiate between the prompt and output, repeats both parts blockwise, and reshapes the SFT mask, as illustrated in Figure 4(b). Based on a more regular attention mask, we combines it with FlexAttention, cutting post-training latency dramatically.
@@ 4.2 Training and Inference Integration
Existing RL efforts in dLLMs (Wang et al., 2025b), as noted earlier, lack inference-engine support and co-design optimization. Concretely, Figure 5(a) illustrates the online-learning loop in existing dLLM-oriented RL. Each training step fetches new data, loads the previous checkpoint for inference, conducts rollout, loads the checkpoint twice for training, trains, and then saves the updated checkpoint back to the file system for reloading in the next step, resulting in an evident waste of IO interactions. Moreover, in earlier works (Zhao et al., 2025a; Zhu et al., 2025b), without the support of an inference engine, the inference of dLLM is slow, further amplifying the overall training inefficiency.
Compared with existing approaches, DiRL tightly couples training and inference as shown in Figure 5(b). We first deploy the model in an API server through LMDeploy (InternLM, 2023), which is the only loading operation in the whole training process. Leveraging LMDeploy’s in-place parameter-update API, we immediately push each training-step checkpoint into the server, eliminating IO between the file system while keeping the API server alive. As shown in Figure 7, these refinements cut per-step latency sharply and dramatically boost the efficiency of dLLM-oriented RL.
@@ 4.4 Main Results
We evaluate our DiRL-8B-Instruct with other baselines on five representative math tasks, MATH500 (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), AIME2024 (MAA, 2024), AIME2025 (MAA, 2025), OlympiadBench (He et al., 2024). Baselines include the blockwise dLLMs SDAR-8B-Chat (Cheng et al., 2025) and TraDo-8B-Instruct (Wang et al., 2025b). We report results for both static and dynamic decoding. For the latter, we set a threshold of 0.9, decoding tokens whose top-1 probability exceeds 0.9 directly. As an external reference, we also list the widely used autoregressive LLM Qwen2.5 Series (Qwen et al., 2024). The results are shown in Table 1.
@@
Table 1: Comprehensive benchmark results of our DiRL-8B-Instruct compared with current dLLMs, SDAR-8B-Chat (Cheng et al., 2025) and TraDo-8B-Instruct (Wang et al., 2025b), as well as well-acknowledged AR model Qwen2.5 Series (Qwen et al., 2024). Each cell presents the accuracy, with best values in bold and suboptimal values underlined, as well as the average number of decoding tokens per step and the average output lengths.
@@ 5.1 Ablation Study
To investigate the sensitivity of the models to hyperparameters within the dynamic decoding strategy, we perform an ablation study on the dynamic sampling threshold $\tau$, adjusting it within the range of $0.5$ to $0.99$, and record the accuracy and average scores of DiRL-8B-Instruct, TraDo-8B-Instruct (Wang et al., 2025b), and SDAR-8B-Chat (Cheng et al., 2025) on five mathematical benchmarks (as illustrated in Figure 8).
@@ 5.2 Future work
Despite the clear gain in training efficiency and the current best-result at 8B scale, several improvements remain for our future work. First, we will scale the approach to larger dLLMs to pursue even stronger performance. Second, although our 8k inference length is already the longest reported for dLLMs, it is still short compared with AR models (Guo et al., 2025; OpenAI, 2024). We have not yet explored test-time scaling or long CoT techniques that could raise the ceiling of model itself (Liu et al., 2025a; Zeng et al., 2024; 2025b). Third, we will incorporate strategies such as dynamic packing (Bai et al., 2024) from AR long-context training to further accelerate our framework. Finally, though currently focused on math tasks, we will extend evaluation to agentic and code tasks (Zeng et al., 2025a), striving to match or surpass AR models across a broader spectrum.
@@ Appendix A More Related Work of RL in dLLMs
Besides TraceRL (Wang et al., 2025c) targeted on blockwise dLLMs, more works have made contributions to adapting RL to Diffusion LLMs. diffu-GRPO (Zhao et al., 2025b) uses a single forward pass to perform “one-step decoding” on masked completion tokens (with optional random prompt masking), approximating the sequence log-probability as the sum of per-token log-probabilities, which enables low-cost policy optimization but introduces substantial approximation bias. To mitigate this issue, Coupled-GRPO (Gong et al., 2025) constructs a pair of negatively correlated estimation samples by applying complementary masks at (t, T-t), thereby reducing the Monte Carlo variance of token log-probability estimates and stabilizing GRPO updates.
@@
Similarly, when performing DPO alignment for diffusion LLMs using ELBO to approximate log-likelihood—which introduces high variance and unstable training—VRPO (Zhu et al., 2025a) systematically reduces the variance of the preference score by increasing the ELBO sampling budget, allocating the entire budget to different timesteps ($n_{t}=n$, $n_{y_{t}}=1$), and letting the current and reference model share the same batch of ($t$, masked $y_{t}$) samples for antithetic sampling. However, because VRPO (Zhu et al., 2025a) must keep the computation graphs of all Monte Carlo samples in memory at once, it easily leads to GPU memory blow-ups and limits the number of samples that can be used. To resolve this, BGPO (Lin et al., 2025) reformulates VRPO’s objective—“taking the exponential of the average log-likelihood difference over all time samples,” which couples all samples—into a linearly separable lower bound over individual samples, enabling per-sample backpropagation and gradient accumulation and reducing memory usage from $O(n_{t})$ to $O(1)$. Other variance-reduction methods optimize sampling strategies, such as GDPO (Rojas et al., 2025), which fixes the noise levels $t$ to a small set of Gauss–quadrature points and performs one-step parallel denoising on generated answers to approximate the sequence-level ELBO, thereby enabling low-variance sequence-level importance ratios.
@@
However, in the above approaches, random sampling and random masking do not seem to resolve the mismatch between the reinforcement learning objective and the actual reasoning trajectory. TraceRL (Wang et al., 2025c) abandons the “random mask + one-step denoise” scheme for obtaining importance sampling ratios. Instead, it segments the generated sequence into trajectories by step/block, uses a shrinkage parameter s to aggregate neighboring steps, and computes token-level ratios within each step by weighting with the old/new policy probability ratios. It then propagates learning signals step by step along the generation process, while introducing a diffusion-based value model to construct step-level GAE advantages.
@@
In addition, to reduce the “zero-gradient” issue in GRPO, IGPO (Zhao et al., 2025c) uses ground-truth fragments as controllable hints to perform inpainting on fully incorrect samples, generating learnable correct trajectories and stabilizing their influence through entropy filtering, thereby restoring effective gradients for diffusion LLMs under sparse rewards.

### v2

@@ 1 Introduction
Large Diffusion Language Models (dLLMs) have become a hot topic in NLP (nie2025large; mercury). The emergence of dLLMs such as LLaDA (nie2025large), Dream (ye2025dream), Mercury (mercury) and Gemini Diffusion (genimid), together with blockwise hybrids like SDAR (cheng2025sdar) that combine diffusion with traditional auto-regression (AR), confirms the scalability of this paradigm (nie2024scaling; gong2024scaling; ni2025training). Building on these results, a growing body of work now seeks to advance their performance in multi-modality (yang2025mmada; you2025llada), long-context modeling (liu2025longllada; he2025ultrallada) and inference efficiency (wu2025fast; wu2025fast2; song2025sparse). Although pre-training of dLLMs is now proven feasible, post-training of dLLMs, especially reinforcement learning (RL), remains underdeveloped, limiting dLLMs’ performance on math tasks and real-world deployment.
@@
The difficulty of dLLM post-training, especially RL, lies in the fact that the logits and derived policy cannot be computed exactly (zhao2025d1; zhu2025llada). In original fully bidirectional dLLMs, the generation order is unconstrained, making teacher-forcing style logit acquisition during SFT infeasible. Injecting uniform random noise into the output fails to reproduce the realistic inference step map, resulting in biased logits and a large mismatch between training and inference objectives (zhao2025d1; wang2025revolutionizing). Furthermore, in the RL stage, the absence of a KV cache further increases computational overhead (liudllm; ma2025dkv; song2025sparse). Most existing dLLM-based RL efforts lack an inference-engine backend, efficient training–inference co-design, and fast rollouts with online model updates, which prevents the practical adoption of mature RL algorithms such as GRPO (deepseekv3; guo2025deepseek). Blockwise dLLMs partially alleviate these issues by restricting generation within blocks, enabling exact logit computation through blockwise forward passes (cheng2025sdar; wang2025revolutionizing). However, they do not fully resolve the train–inference mismatch in post-training, nor do they address the efficiency and algorithmic challenges of dLLM RL. How to achieve consistent training and inference while enabling scalable RL for dLLMs remains underexplored. To fill this gap, we introduce our dLLM RL algorithm DiPO, together with the training framework DiRL, which enforces training–inference consistency and enables efficient rollouts and policy optimization for blockwise dLLMs. We further present the state-of-the-art dLLM DiRL-8B-Instruct, as shown in Figure 2, Figure 2 and Figure 3.
@@
At the algorithm level, DiPO leverages the good property of blockwise dLLM to achieve the first unbiased GRPO implementation for dLLMs through efficient, unbiased logit computation. At the framework level, DiRL supports the two-stage (SFT-RL) post-training of dLLMs, aligning training and inference objectives while surpassing existing methods in efficiency as shown in Figure 2. Concretely, we integrate the efficient inference property of blockwise dLLMs with the efficient FlexAttention interface (dong2024flex) and LMDeploy framework (2023lmdeploy) to enable fast rollout and online model updates in the API server. At the model level, based on high-quality math datasets, we train DiRL-8B-Instruct from SDAR-8B-Chat and achieve best performance on math tasks in dLLMs, even outperform Qwen2.5 Series (qwen2024qwen25technicalreport), the widely-acknowledged larger AR model, in AIME24, AIME25 (aime2024; aime2025) and OlympiadBench (he2024olympiadbench) as shown in Figure 2. Our contributions can be summarized as follows.
@@ 2.1 Block Diffusion Language Model
Blockwise Diffusion Language Models, as exemplified by BD3-LMs (arriola2025block) and SDAR (cheng2025sdar), primarily unify the global sequential dependency of AR models with the local parallel generation capability of dLLMs through a Semi-Autoregressive generation paradigm.
@@
Thanks to the combination of intra-block AR and inner-block diffusion, the KV cache mechanism can be applied to blockwise dLLMs, thus enhancing the computational efficiency of traditional dLLMs and simplifying the likelihood computation (cheng2025sdar; wang2025diffusion).
@@ 2.2 RL and its application in dLLM
Thanks to the introduction of blockwise dLLMs (cheng2025sdar; wang2025diffusion), the computational overhead of log-probability is greatly reduced. Building on this, wang2025revolutionizing proposes TraceRL, which leverages the algorithmic efficiency of blockwise dLLMs to enable faster unbiased logit computation. However, due to the lack of attention acceleration during training and the absence of tight integration between training and inference frameworks, TraceRL still leaves significant room for engineering optimization. Moreover, it does not implement more fine-grained RL algorithms such as GRPO. To address these limitations, we introduce our efficient training framework DiRL and DiPO, the first unbiased implementation of GRPO for dLLMs.
@@ 3.1 SFT stage
We train on the OpenR1-Math dataset (openr1) distilled from GLM-4.6 (zeng2025glm; glm46). We chose GLM-4.6 because currently, it is almost the best-performing open-source non-reasoning LLM on math tasks and yields reasoning trajectories of manageable length. As long-reasoning evaluation for diffusion models is still scarce, we cap the reasoning length at 8k token length, which is already the longest inference length reported for dLLMs.
@@ 3.2 RL stage
Moreover, we integrate the DAPO(yu2025dapoopensourcellmreinforcement) algorithm, which uses the token-level policy gradient:
@@
We train on the Big-Math dataset (albalak2025big), a large-scale, high-quality math dataset verfied by math-verify, and specially designed for RL.
@@ 4 Engineering Optimization
As shown in Figure 2, the DiRL framework incorporates efficiency designs at both the model and system levels. At the model level, blockwise dLLM (cheng2025sdar) with FlexAttention (dong2024flex) compute unbiased logits efficiently, and at the framework level, tight integration of the LMDeploy inference engine (2023lmdeploy) with the open-source distributed training library LLaMA-Factory (zheng2024llamafactory) enables online model updates for highly efficient training.
@@ 4.1 Blockwise dLLM with FlexAttention
At the model level, as mentioned earlier, we use blockwise dLLM, which is more efficient for logic computation than the original dLLMs in post-training. This is because it maintains an AR paradigm between blocks, performing denoising only within each block. Thanks to this design, as Figure 4(a) shows, TraceRL (wang2025revolutionizing) can parallelize denoising across blocks by repeating the output twice and altering the attention mask, thus enabling faster SFT training.
However, the dominant FlashAttention interface (dao2022flashattention; dao2023flashattention) cannot handle the complex attention masks required by blocksize dLLMs, limiting the practical training throughput. Fortunately, PyTorch’s FlexAttention (dong2024flex) fills this gap with an efficient operator that accepts fine-grained masks. DiRL does not differentiate between the prompt and output, repeats both parts blockwise, and reshapes the SFT mask, as illustrated in Figure 4(b). Based on a more regular attention mask, we combines it with FlexAttention, cutting post-training latency dramatically.
@@ 4.2 Training and Inference Integration
Existing RL efforts in dLLMs (wang2025revolutionizing), as noted earlier, lack inference-engine support and co-design optimization. Concretely, Figure 5(a) illustrates the online-learning loop in existing dLLM-oriented RL. Each training step fetches new data, loads the previous checkpoint for inference, conducts rollout, loads the checkpoint twice for training, trains, and then saves the updated checkpoint back to the file system for reloading in the next step, resulting in an evident waste of IO interactions. Moreover, in earlier works (zhao2025d1; zhu2025llada), without the support of an inference engine, the inference of dLLM is slow, further amplifying the overall training inefficiency.
Compared with existing approaches, DiRL tightly couples training and inference as shown in Figure 5(b). We first deploy the model in an API server through LMDeploy (2023lmdeploy), which is the only loading operation in the whole training process. Leveraging LMDeploy’s in-place parameter-update API, we immediately push each training-step checkpoint into the server, eliminating IO between the file system while keeping the API server alive. As shown in Figure 7, these refinements cut per-step latency sharply and dramatically boost the efficiency of dLLM-oriented RL.
@@ 4.4 Main Results
We evaluate our DiRL-8B-Instruct with other baselines on five representative math tasks, MATH500 (hendrycks2021math), GSM8k (cobbe2021gsm8k), AIME2024 (aime2024), AIME2025 (aime2025), OlympiadBench (he2024olympiadbench). Baselines include the blockwise dLLMs SDAR-8B-Chat (cheng2025sdar) and TraDo-8B-Instruct (wang2025revolutionizing). We report results for both static and dynamic decoding. For the latter, we set a threshold of 0.9, decoding tokens whose top-1 probability exceeds 0.9 directly. As an external reference, we also list the widely used autoregressive LLM Qwen2.5 Series (qwen2024qwen25technicalreport). The results are shown in Table 1.
@@
Table 1: Comprehensive benchmark results of our DiRL-8B-Instruct compared with current dLLMs, SDAR-8B-Chat (cheng2025sdar) and TraDo-8B-Instruct (wang2025revolutionizing), as well as well-acknowledged AR model Qwen2.5 Series (qwen2024qwen25technicalreport). Each cell presents the accuracy, with best values in bold and suboptimal values underlined, as well as the average number of decoding tokens per step and the average output lengths.
@@ 5.1 Ablation Study
To investigate the sensitivity of the models to hyperparameters within the dynamic decoding strategy, we perform an ablation study on the dynamic sampling threshold $\tau$, adjusting it within the range of $0.5$ to $0.99$, and record the accuracy and average scores of DiRL-8B-Instruct, TraDo-8B-Instruct (wang2025revolutionizing), and SDAR-8B-Chat (cheng2025sdar) on five mathematical benchmarks (as illustrated in Figure 8).
@@ 5.2 Future work
Despite the clear gain in training efficiency and the current best-result at 8B scale, several improvements remain for our future work. First, we will scale the approach to larger dLLMs to pursue even stronger performance. Second, although our 8k inference length is already the longest reported for dLLMs, it is still short compared with AR models (guo2025deepseek; OpenAI2024o1). We have not yet explored test-time scaling or long CoT techniques that could raise the ceiling of model itself (liu2025thus; zeng2024scaling; zeng2025revisiting). Third, we will incorporate strategies such as dynamic packing (bai2024longalign) from AR long-context training to further accelerate our framework. Finally, though currently focused on math tasks, we will extend evaluation to agentic and code tasks (zeng2025glm), striving to match or surpass AR models across a broader spectrum.
@@ Appendix A More Related Work of RL in dLLMs
Besides TraceRL (wang2025revolutionizingreinforcementlearningframework) targeted on blockwise dLLMs, more works have made contributions to adapting RL to Diffusion LLMs. diffu-GRPO (zhao2025d1scalingreasoningdiffusion) uses a single forward pass to perform “one-step decoding” on masked completion tokens (with optional random prompt masking), approximating the sequence log-probability as the sum of per-token log-probabilities, which enables low-cost policy optimization but introduces substantial approximation bias. To mitigate this issue, Coupled-GRPO (gong2025diffucoderunderstandingimprovingmasked) constructs a pair of negatively correlated estimation samples by applying complementary masks at (t, T-t), thereby reducing the Monte Carlo variance of token log-probability estimates and stabilizing GRPO updates.
@@
Similarly, when performing DPO alignment for diffusion LLMs using ELBO to approximate log-likelihood—which introduces high variance and unstable training—VRPO (zhu2025vrporethinkingvaluemodeling) systematically reduces the variance of the preference score by increasing the ELBO sampling budget, allocating the entire budget to different timesteps ($n_{t}=n$, $n_{y_{t}}=1$), and letting the current and reference model share the same batch of ($t$, masked $y_{t}$) samples for antithetic sampling. However, because VRPO (zhu2025vrporethinkingvaluemodeling) must keep the computation graphs of all Monte Carlo samples in memory at once, it easily leads to GPU memory blow-ups and limits the number of samples that can be used. To resolve this, BGPO (lin2025boundaryguidedpolicyoptimizationmemoryefficient) reformulates VRPO’s objective—“taking the exponential of the average log-likelihood difference over all time samples,” which couples all samples—into a linearly separable lower bound over individual samples, enabling per-sample backpropagation and gradient accumulation and reducing memory usage from $O(n_{t})$ to $O(1)$. Other variance-reduction methods optimize sampling strategies, such as GDPO (rojas2025improvingreasoningdiffusionlanguage), which fixes the noise levels $t$ to a small set of Gauss–quadrature points and performs one-step parallel denoising on generated answers to approximate the sequence-level ELBO, thereby enabling low-variance sequence-level importance ratios.
@@
However, in the above approaches, random sampling and random masking do not seem to resolve the mismatch between the reinforcement learning objective and the actual reasoning trajectory. TraceRL (wang2025revolutionizingreinforcementlearningframework) abandons the “random mask + one-step denoise” scheme for obtaining importance sampling ratios. Instead, it segments the generated sequence into trajectories by step/block, uses a shrinkage parameter s to aggregate neighboring steps, and computes token-level ratios within each step by weighting with the old/new policy probability ratios. It then propagates learning signals step by step along the generation process, while introducing a diffusion-based value model to construct step-level GAE advantages.
@@
In addition, to reduce the “zero-gradient” issue in GRPO, IGPO (zhao2025inpaintingguidedpolicyoptimizationdiffusion) uses ground-truth fragments as controllable hints to perform inpainting on fully incorrect samples, generating learnable correct trajectories and stabilizing their influence through entropy filtering, thereby restoring effective gradients for diffusion LLMs under sparse rewards.